# Understanding intra-node communication in HPC systems and Datacenters

### v1

@@ 4.1 Simulation model validation
Note that the intra-node network model is fully configurable to behave not only as PCIe but also to be as generic as possible. In the following sections, we extend the analysis to the results of scale-out simulations using generic intra- and inter-node configurations.
@@ 4.2 Scale-out simulation experiments
To analyze the impact of intra-node communication involving multiple accelerators per node on large inter-node networks, we have configured the simulation model (see Section 3.3) to operate similarly to those intra- and inter-node networks commonly used in Datacenters and Supercomputers. This section shows the simulation results for large network configurations (see Table 3) when different realistic traffic patterns are used. In the following sections, we describe the experiments’ configuration (i.e., the intra- and inter-node network configuration and the communication patterns) and analyze the simulation results.
4.2.1 Intra- and inter-node network configurations
We have modeled the end-node architecture similarly to that of modern compute nodes. We assume each end node includes $8$ accelerators interconnected by a high-speed intra-node network. Each accelerator has a NIC to enable data generation in the intra-node network. We assume that the intra-node network can operate at the following aggregated speeds:
  • $128$ GB/s: Each accelerator NIC has a data rate of $128$ Gbps, so the aggregated intra-node network bandwidth among all accelerators is $8\times 128=1024$ Gbps (i.e., $128$ GB/s).
  • $256$ GB/s: Each accelerator NIC data rate is $256$ Gbps, the aggregated intra-node network bandwidth being $256$ GB/s.
  • $512$ GB/s: Each accelerator NIC data rate is $512$ Gbps, the aggregated intra-node network bandwidth being $512$ GB/s.
These link speeds and aggregated bandwidth are consistent with several intra-node network technologies, such as PCIe or Infinity Fabric [18], which use a bandwidth per link similar to the inter-node network link. Unfortunately, the aggregated bandwidth of the intra-node network is much larger than that of a single inter-node network link, so this shows clearly that end-node NICs might be the bottleneck if all the end-node devices would want to communicate at full speed to devices at different end nodes.
Table 3 shows the inter-node network configurations. We have configured $32$- and $128$-node inter-node networks using a Real-Life Fat-Tree (RLFT) topology [19]. As we assume $8$ accelerators per node, the total number of accelerators interconnected using the intra- and inter-node networks is $256$ and $1024$, respectively.
Table 3: Network configurations used in the experiments.
@@ 4.2.1 Intra- and inter-node network configurations
# Nodes Accelerators Topology Inter-node switches Routing
1 32 256 RLFT 12 D-mod-K
2 128 1024 RLFT 24 D-mod-K
We assume $400$ Gbps links for the inter-node network, consistent with current high-speed networking technologies such as InfiniBand or Gigabit Ethernet. The latency for the first flit of a packet traversing the inter-node network is set to $6$ ns. The switching policy is virtual cut-through, and the flow control mechanism is credit-based. We assume the $D$-mod-$K$ deterministic routing algorithm [20] for both network configurations, which efficiently balances traffic flows paths among available shortest routes.
Regarding the performance metrics, we measured the intra-node latency that measures the time required for an intra-node packet to be sent and received within the same end node. It excludes packets that traverse the inter-node network. We also modeled the intra-node throughput that measures the average data rate per time unit (in GB/s) transmitted within the intra-node networks. We have also measured the inter-node throughput that measures the data rate transmitted through the inter-node network. Finally, we also measure the Flow Completion Time (FCT) that measures the total time taken for a message to be transmitted from its generation at the source accelerator until its reception at the destination.
4.2.2 Communication patterns configuration
Our experiments have evaluated the impact of intra-node and inter-node communication on system performance. To achieve this, we explore different DP, PP, and PP configurations (see Section 3.4). We have considered five communication patterns: $C1$, $C2$, $C3$, $C4$, $C5$, which differ on the amount of traffic generated at the accelerators addressed to the intra- and inter-node network. In other words, we vary the amount of messages generated due to data and model parallelism. However, this configuration does not imply that these communication patterns belong to specific LLM models. Note that determining the most efficient and optimal configuration for the different types of parallelism (e.g., DP, PP or TP) for a specific LLM is beyond the scope of our work. For such purposes, existing tools such as Calculon [21] can provide valuable insights.
@@ 4.2.2 Communication patterns configuration
4.2.3 Analysis of simulation results
(a) Throughput - 128 GB/s intra-node.
(b) Throughput - 256 GB/s intra-node.
(c) Throughput - 512 GB/s intra-node.
(d) Latency - 128 GB/s intra-node.
(e) Latency - 256 GB/s intra-node.
(f) Latency - 512 GB/s intra-node.
Figure 5: Intra-node network performance vs Traffic Load (%) in a 32-node RLFT topology (network config. #1 in Table 3).
(a) Throughput - 128 GB/s intra-node.
(b) Throughput - 256 GB/s intra-node.
(c) Throughput - 512 GB/s intra-node.
(d) FCT - 128 GB/s intra-node.
(e) FCT - 256 GB/s intra-node.
(f) FCT - 512 GB/s intra-node.
Figure 6: Inter-node network performance vs Traffic Load (%) in a $32$-node RLFT topology (network config. #1 in Table 3).
(a) Throughput - 128 GB/s intra-node.
(b) Throughput - 256 GB/s intra-node.
(c) Throughput - 512 GB/s intra-node.
(d) Latency - 128 GB/s intra-node.
(e) Latency - 256 GB/s intra-node.
(f) Latency - 512 GB/s intra-node.
Figure 7: Intra-node network performance vs Traffic Load (%) for a $128$-node RLFT topology (network config. #1 in Table 3).
(a) Throughput on network of 128 GB/s.
(b) Throughput on network of 256 GB/s.
(c) Throughput on network of 512 GB/s.
(d) FCT on network of 128 GB/s.
(e) FCT on network of 256 GB/s.
(f) FCT on network of 512 GB/s.
Figure 8: Inter-node network performance vs Traffic Load (%) for a $128$-node RLFT topology (network config. #1 in Table 3).
Figure 5 illustrates the intra- and inter-node performance results versus traffic load for a $32$-node RLFT network (i.e., $256$ accelerators) when the $C1$-to-$C5$ traffic patterns are generated into the three different intra-node configurations. Note that inter-node network NICs can handle up to $400$ Gbps (i.e., $50$ $\mathrm{GB}$/s), while the intra-node network can inject up to $128$, $256$, or $512$ $\mathrm{GB}$/s, depending on the network configuration. As we can see in Figures 5(a), 5(b) and 5(c) the maximum aggregated intra-node network throughput increases proportionally to the aggregated intra-node bandwidth. The saturation point [^2] Note that the network saturation point is where throughput drops to zero due to NIC buffers collapse and packets are not able to reach the destination during the simulation time, and latency skyrockets. in terms of generated intra-node traffic load is reached when more inter-node traffic is generated (e.g., $C1$), and this situation becomes worse as the aggregated intra-node bandwidth increases (see Figure 5(c)). The same trends as for intra-node throughput can be observed if we look at intra-node latency metrics (see Figures 5(d), 5(e) and 5(f)).
Figure 6 shows the inter-node network throughput and FCT metrics for the same scenarios. As we can see, communication patterns $C1$ and $C2$ achieve the best inter-node throughput even though they saturate soon in the intra-node network. Presumably, this is because these communication patterns generate the largest amount of inter-node traffic. By contrast, as it can be seen in Figures  6(d), 6(e) and 6(f), where we can see that communication patterns $C1$ and $C2$ achieve the worst performance due to the intra-node bottleneck described above. Another interesting aspect regarding latencies (intra-node latency and FCT) is that they increase exponentially as the inter-node network injects traffic. This indicates the presence of two distinct bottlenecks: at the source end-node and the destination end-node. At both sides, messages destined for the inter-node network must traverse their respective NICs, which are bottlenecked due to the saturation. The source end-node’s NIC is limited to handling a maximum of $50$ $\mathrm{GB}$/s. Meanwhile, the destination end-node’s NIC packetizes inter-node packets (of $4$ $\mathrm{KiB}$) into intra-node packets of $128$ $\mathrm{B}$, injecting a large number of packets from the inter-node network into the intra-node network.
Focusing on traffic patterns, as TP increases, leading to greater traffic injection into the inter-node network, the overall system performance degrades due to increased contention at the end-node NICs and the emergence of bottlenecks. This does not necessarily indicate a reduction in LLM computational performance, as multiple accelerators still collaborate to accelerate sub-layer processing. However, it is crucial to analyze network impact, as excessive communication delays can degrade overall LLM performance.
Figures 5 and 6 show that increasing intra-node network aggregated bandwidth may negatively affect overall network performance under certain communication patterns (e.g., LLMs with a high degree of TP). Note that increasing intra-node bandwidth enhances performance when traffic remains confined within the intra-node network (as in configuration $C5$), but this differs for other configurations ($C1$, $C2$, $C3$, and $C4$), where the additional traffic injected into the inter-node network spoils intra-node network throughput as intra-node bandwidth increases. Conversely, the inter-node network benefits from this increased intra-node bandwidth, as more traffic can be handled.
Figures 7 and 8 show that increasing the network size to a $128$-node RLFT topology (i.e., $1024$ accelerators) does not alter the trends observed for the $32$-node configuration. For the intra-node network, the throughput and latency trends remain consistent. The throughput scales proportionally with the number of nodes; as there are four times more accelerators in the $128$-node topology. However, the intra-node latency remains unchanged because the intra-node network configuration and the number of accelerators per node are identical to those in the smaller topology. Thus, similar observations can be made regarding the inter-node network and FCT. Since the bottleneck originates in the intra-node network, the inter-node network exhibits the same tendencies observed in the smaller topology. Throughput in the inter-node network increases proportionally with the number of nodes, reflecting the higher data exchange between nodes. In summary, these results demonstrate the impact of the intra-node network traffic on the overall system performance. Bottlenecks within the intra-node network have significant downstream effects on the inter-node network, underscoring the importance of optimizing intra-node communication for scaling large systems effectively.

### v2

@@ 4.1 Simulation model validation
Note that the intra-node network model is fully configurable to behave not only as PCIe but also to be as generic as possible. In the following sections, we extend the evaluation to scale-up and scale-out simulations using generic but realistic intra- and inter-node configurations.
4.2 Experiments configurations
In the simulator, we have configured the end-node architecture and the intra- and inter-node networks similarly to modern Supercomputers and Data Centers composed of heterogeneous end nodes. At each end node, we assume up to $8$ accelerators (e.g., GPUs) interconnected by a high-speed intra-node network. Each accelerator has a single NIC to enable data generation to the intra-node network. Also, a NIC connects the intra-node network to the inter-node network. We assume three configurations for the intra-node network using different speeds at accelerator NICs:
  • Configuration #1 - $128$ Gbps/accelerator. Each accelerator NIC has a maximum generation rate of $128$ Gbps, so the aggregated intra-node network bandwidth among all accelerators ranges from $128$ (11 accelerator) to $1024$ Gbps ($8$ accelerators).
  • Configuration #2 - $256$ Gbps/accelerator. Each accelerator NIC can generate traffic at a maximum rate of $256$ Gbps, so the aggregated intra-node network bandwidth varies from $256$ (11 accelerator) to $2048$ Gbps ($8$ accelerators).
  • Configuration #3 - $512$ Gbps/accelerator. Each accelerator NIC can generate traffic at a maximum rate of $512$ Gbps, so the aggregated intra-node network bandwidth varies from $512$ (11 accelerator) to $4096$ Gbps ($8$ accelerators).
These link speeds and aggregated bandwidth are consistent with several intra-node network technologies, such as PCIe or Infinity Fabric [18], which utilize a bandwidth per link similar to that of the inter-node network link. Note that the aggregated bandwidth of the intra-node network can be much larger than that of a single inter-node network NIC, which is assumed to be $400$ Gbps (i.e., 50 GB/s) according to current commercial speeds, such as InfiniBand or Gigabit Ethernet. This clearly shows that end-node NICs can be the bottleneck if all the end-node devices want to communicate at full speed with devices at different end nodes.
To measure the impact of a varying number of accelerators in generating bottlenecks, we assume $1$, 22, 44, or $8$ accelerators per end node connected to a single intra-node network switch. That switch is also connected to the end-node NIC via an intra-node link. At this switch, we assume buffering only at input ports of size 128 $\mathrm{KiB}$ (i.e., input queued IQ architecture), virtual cut-through (VCT) switching, and credit-based flow control. Additionally, switch arbitration is made utilizing the iSlip algorithm [19], so that each switch output port independently selects which crossing request of packets at input-port buffers is chosen, allowing that packet to cross through that output port. This decision is based on a round-robin (RR) policy among all the input ports or choosing the oldest crossing request (aging). Note that the number of ports of this intra-node switch depends on the number of accelerators per end node. The link bandwidth of each switch port is identical to that of the accelerator NIC.
The inter-node network architecture is similar to that of the intra-node network. Likewise, we have configured high-speed IQ switches of size 128 $\mathrm{KiB}$ using VCT, credit-based flow control, and RR- or Age-based iSlip scheduling. We have selected a Real-Life Fat-Tree (RLFT) network topology [20], widely used in current Supercomputers and data centers. We assume the $D$-mod-$K$ deterministic routing algorithm [21], which efficiently balances traffic flow paths among available shortest-path routes in the network, and achieves similar performance to adaptive routing [22]. Finally, regarding the link flying time of packets, we assume that the latency for the first flit of any packet traversing the inter-node network is set to $25$ ns/meter.
Regarding the performance metrics, we have modeled the intra-node throughput that measures the average data rate per time unit (in Gbps) of the traffic that is only exchanged among accelerators within the intra-node network. We also measure the inter-node throughput, the data rate transmitted through the inter-node network. These two metrics will allow us to analyze what happens in each scenario. Regarding latency, we decompose the latency for a complete message (i.e., before packetization) into several components and measure different types of latency. Figure 8 represents the seven latency metrics we measured.
Figure 8: Latency metrics experienced by a traffic flow traversing the intra-node (red flow) and inter-node (green flow) networks.
This Figure shows a red flow from $Acc_{0}$ of $Server_{1}$ to $Acc_{N}$ of $Server_{M}$, which illustrates inter-node communication. Also, we show a green flow from $Acc_{0}$ to $Acc_{N}$ within $Server_{1}$, illustrating intra-node communication. First, we can observe the Source Accelerator Latency, which indicates the latency from the creation of the packet to its transmission. The Intra-node Network Latency models the time since a packet leaves the Source Acc. until it reaches the end-node NIC. At the Source end node, we finally have the Source NIC Latency, which indicates the time required to receive enough intra-node packets to be aggregated into an inter-node packet MTU. Then, it indicates the time between the arrival of the first intra-node packet and the sending of the inter-node MTU. At the destination end node, there are the same latencies, but in the opposite direction. Finally, we can observe the Inter-node Network Latency, which represents the time needed to travel through the inter-node network.
Regarding the traffic patterns, we want to evaluate the impact of intra- and inter-node communication on the overall network performance. We assume the traffic model described in Section 3.4 based on LLMs training, which combines several configurations for the data parallelism (DP), Tensor Parallelism (TP), and Pipeline Parallelism (PP) into five communication patterns: $C1$, $C2$, $C3$, $C4$, $C5$. They differ in the traffic rate generated at the accelerators addressed to the intra- and inter-node network. Note that this configuration does not imply that these communication patterns belong to any specific LLM model. Modeling specific LLMs’ traffic patterns, the different types of parallelism (e.g., DP, PP, or TP), or determining the most efficient and optimal configuration for those LLMs is beyond the scope of this paper. For such purposes, existing tools such as Calculon [23] can provide valuable insights, but this analysis and modeling effort is left for future work.
[…]
4.3 Intra-node overhead analysis
As described in Section 3.5, there is an overhead produced when messages generated at a given accelerator are split into packets in the intra-node network, then recomposed again into larger inter-node packets, then split again into intra-node packets at the receiving end node, and finally received in the remote accelerator. In our experiments, we assume intra-node packets with a header size of 20 $\mathrm{B}$, and a payload of 128 $\mathrm{B}$, and inter-node packets whose header is 64 $\mathrm{B}$ and the payload is $4032$ $\mathrm{B}$. Substituting these values into Equation 2, we obtain the following traffic overhead factor:
$\text{TrafficOverhead}=\frac{\frac{20}{128}+1}{\frac{64}{4032}+1}=\frac{1.15625}{1.01587}\approx 1.14$
By applying this overhead to Equation 3, where we have adjusted the Model_adjustment field to $15703.9$ as provided by the mathematical analysis made as described in Section 3.5), we can express the formula as follows:
$\text{Throughput (Gbps)}=\frac{$15703.9$}{1.14\times\text{TrafficInterNode (\%)}}\times\text{NumNodes}$
This formula can be generalized for any number of nodes and used to estimate the network performance bound, assuming the overhead introduced by the conversion of packets across intra- and inter-node networks. This relationship remains unchanged regardless of the intra-node network speed, as the header and payload sizes are fixed.
4.4 Scale-Up simulation results
Next, we evaluate the experiment results as the number of accelerators per end node increases while keeping the number of nodes in the system. We have configured a $32$-node system interconnected using a $2$-stage RLFT topology comprising twelve $8$-port switches. We increase the number of accelerators per server node ($1$, 22, 44, and $8$), so the total number in the system ranges from $32$ to $256$. We configure the rest of the inter- and intra-node network parameters as described in Section 4.2. We have also configured different speeds for the accelerator NICs (e.g., $128$, $256$, and $512$ Gbps). Table 2 summarizes the intra-node configurations.
Table 2: Scale-Up network configurations.
# Intra-node link speed Total # of Accelerators
1 $128$ Gbps/accelerator $32$, $64$, $128$, $256$
2 $256$ Gbps/accelerator $32$, $64$, $128$, $256$
3 $512$ Gbps/accelerator $32$, $64$, $128$, $256$
4.4.1 Analysis of switch arbiter policies
First, we analyze whether the switch arbitration policy impacts overall system performance. We compare two arbitration schemes: Round-Robin (RR) and Age-Based for all the intra- and inter-node switches. We assume the intra-node network configurations #1, #2, and #3 from Table 2 only using $8$ accelerators per node, so there are $256$ accelerators in the system. Figure 9 shows the average network throughput in the intra-node network versus the traffic load (between 0% and 100%) generated at accelerators’ NICs, when the five traffic patterns C1, C2, C3, C4, and C5 are generated. As we can see, the performance differences between the two arbitration schemes are negligible across all configurations. This indicates that the arbitration policy does not significantly affect the system throughput in the assumed scenarios.
(a) Age Based - conf #1.
(b) Age Based - conf #2.
(c) Age Based - conf #3.
(d) Round Robin - conf #1.
(e) Round Robin - conf #2.
(f) Round Robin - conf #3.
Figure 9: Intra-node network performance with different arbiters as a function of traffic load (%) in a 32-node RLFT topology.
It is also worth mentioning that, as expected, when the intra-node link speed augments, the throughput saturation point is reached with a smaller aggregated traffic load for those traffic patterns generating a higher percentage of inter-node traffic. For instance, the traffic pattern C1, which generates 80% of intra- and 20% of inter-node traffic, saturates the network with 60% of aggregated generation load in network configuration #1 (see Figures 9(a) and 9(d)), with 30% of traffic load for network configuration #2, and with 10% of traffic load for network configuration #3.
Figure 10 shows the aggregated latency results for the scenarios mentioned above. Each vertical bar is divided into seven sections for the seven latency components described in Figure 8. The “X” axis shows ten load instants (from 0% to 100%) and one bar per traffic pattern C1-C4. Note that the C5 traffic pattern has not been shown since there is no inter-node communication. These barplots also show that most latency is accumulated in the source intra-node network when saturation happens (red color). When the intra-node network speed augments (network configurations #2 and #3), the source NIC, inter-node network, and destination NIC latency increase because the amount of traffic the accelerators send through the inter-node network also augments.
(a) Age Based - conf #1.
(b) Round Robin - conf #1.
(c) Age Based - conf #2.
(d) Round Robin - conf #2.
(e) Age Based - conf #3.
(f) Round Robin - conf #3.
Figure 10: Packet Latency divided into components versus traffic load for the traffic patterns C1-C4 and network configurations #1, #2, and #3 in Table 2.
Regarding the arbitration policies, there are negligible differences between Age-based and Round Robin policies in terms of network throughput, while there are some variations in latency behavior, which do not significantly impact system performance. We can see that Age-based arbitration scenarios experience 20% more aggregated latency under saturation compared to Round Robin policies, since the former prioritizes older crossing requests within network switches. Another important aspect, later analyzed, is the biggest portion of the aggregated latency, which is consumed at the Source intra-node network under saturation. But this portion is shared with Inter-node network latency and Source and Destination NIC when the accelerators NIC speed increases (see configurations #2 and #3). Since the Round Robin arbitration is commonly used and easy to implement in hardware, we have selected this policy for rest of the experiments of the paper hereafter.
4.4.2 Analysis of the intra-node network speed
This section analyzes the intra- and inter-node network performance when the intra-node network scales up in terms of the number of accelerators per end node and the intra-node link speed. Figure 11 illustrates the intra- and inter-node network throughput of a $32$-node RLFT topology, when C1 to C5 traffic patterns are generated. We assume that the number of accelerators per end node is $1$, 22, 44, or $8$, so that the total number of accelerators is $32$, $64$, $128$, and $256$ for the entire system. We also assume that each accelerator can generate up to $128$ Gbps of traffic to the intra-node network.
(a) 11 Acc./Node.
(b) 22 Acc./Node.
(c) 44 Acc./Node.
(d) 88 Acc./Node.
Figure 11: Network throughput as a function of traffic load (%) in a 32-node RLFT topology (network configuration #1 in Table 2). Total network throughput is shown in the top figures, while inter-node throughput is depicted in the bottom ones.
Figure 11(a) shows throughput results (i.e., the amount of information received per accelerator) with a single accelerator per node (i.e., $32$ accelerators in total and $128$ Gbps per accelerator). In this case there is no possible intra-node communication among accelerators within a single end node. For instance, for C1 there is a 20% of inter-node traffic, i.e., each accelerator generates $128\times 0.2=25.6$ Gbps, meaning that the total generation rate in the network can be up to $25.6\times 32$accelerators=819.2=819.2 Gbps. Therefore, the throughput assuming no saturation at maximum load (in this case $25.6$ Gbps per accelerator) is $819.2$ Gbps. Note that single-accelerator end nodes are no longer practical, and multi-accelerator end nodes have become essential. For instance, one of the reasons is the increasingly popular use of LLMs [24] and their growing amount of memory requirements, where Tensor Parallelism (TP) is crucial for partitioning models across multiple accelerators. TP is communication-intensive, requiring frequent message exchanges for All-Reduce operations.
Increasing the number of accelerators per end node to $2$ or $4$ increments the network throughput as well. Figures 11(b) and 11(c) show that there is no saturation for server nodes with $2$ and $4$ accelerators per node, regardless of the traffic pattern, reaching the maximum bandwidth of $8192$ Gbps and $16384$ Gbps, respectively. This throughput is achieved with $64$ and $128$ accelerators generating traffic at full speed (i.e., $128$ Gbps) without any bottleneck in the network. The link between the intra-node network and the NIC (see Figure 8) remains unsaturated. Therefore, the intra and inter-node networks achieve the maximum performance for all the traffic pattern configurations (C1-C4), even for C5, which does not involve inter-node traffic.
Figure 11(d) shows the results with $8$ accelerators per server node. For traffic pattern C5 (i.e., 100% intra-node traffic), the inter-node network is unused, so the system achieves its maximum possible throughput, which is $128~\text{Gbps}/\text{accelerator}\times 8~\text{accelerators}/\text{node}\times 32~\text{nodes}=$3276832768Gbps. In traffic configurations C4 and C3 (i.e., 95% and 90% of intra-node traffic, respectively) do not saturate the NIC link. Considering that the maximum traffic an end node can inject into the inter-node network is $128~\text{Gbps}/\text{accelerator}\times 8~\text{accelerators}=$1024$~\text{Gbps}$, each end-node NIC only receives packets at a rate of 51.2 Gbps for traffic pattern C4 (i.e., 5% of inter-node traffic) and 102.4 Gbps for C3 (i.e., 10% of inter-node traffic), respectively. Thus, the link connecting the intra-node switch to the end-node NIC, whose maximum bandwidth is $128$ Gbps, can accommodate the intra-node load that the $8$ accelerators send to the inter-node network without becoming a bottleneck.
By contrast, traffic patterns C2 and C1 (85% and 80% of intra-node traffic, respectively) exceed the intra-node network link maximum capacity (i.e., $128$ Gbps), since $8$ accelerators per end node generate a traffic load of $153.6$ Gbps for C2 (i.e., 15% of inter-node traffic) and $204.8$ Gbps for C1 (i.e., 20% of inter-node traffic), respectively. In these scenarios, the intra-node link becomes a bottleneck. When scaling up the number of accelerators per end node, it is essential to consider the intra-node link bandwidth and the amount of traffic the accelerators can generate to the inter-node network, as excessive traffic routed through the end-node NIC (either incoming or outgoing) can degrade overall system performance. Note that, due to the mentioned bottleneck, the inter-node network cannot reach its theoretical maximum throughput of $400~\text{Gbps}\times 32~\text{nodes}=$ 1280012800Gbps, as the traffic injection rate from the intra-node network is limited by its maximum link bandwidth, i.e., $128$ Gbps. Therefore, for $128$ Gbps intra-node links, the appropriate number of accelerators per end node is $1$, 22, or $4$, while $8$ accelerators saturate the intra-node network when there is a small percentage of traffic generated to the inter-node network.
Figure 12 shows throughput results for network configuration #2 (see Table 2), when intra-node link speed is $256$ Gbps. In this scenario, the saturation point is reached with only $4$ accelerators and traffic pattern C2 (i.e., 15% of inter-node traffic). Considering the maximum bandwidth all the accelerators of an end node can generate, i.e., $256~\text{Gbps}/\text{accelerator}\times 8~\text{accelerators}=$20482048Gbps. Note that the bottleneck between the intra-node network and the NIC is still present. Indeed, system performance degrades more rapidly compared to network configuration #1.
(a) 11 Acc./Node.
(b) 22 Acc./Node
(c) 44 Acc./Node
(d) 88 Acc./Node
Figure 12: Network performance as a function of traffic load (%) in a 32-node RLFT topology (network config. #2 in Table 2). Total network throughput is shown in the top figures, while inter-node throughput is depicted in the bottom ones.
Figure 13 shows throughput results for network configuration #3 (see Table 2). In this scenario, the link between the intra-node network and the NIC can operate at $512$ Gbps, and the inter-node network links operate at $400$ Gbps which is not enough to alleviate the bottleneck between the inter-node switch and the end-node NIC when accelerators exchange a small amount of traffic. Note that the network saturates with just $2$ accelerators per node generating 6060% of traffic load and traffic pattern C1.
(a) 1 Acc./Node
(b) 2 Acc./Node.
(c) 4 Acc./Node.
(d) 8 Acc./Node.
Figure 13: Network performance as a function of traffic load (%) in a 32-node RLFT topology (network configuration #3 from Table 2). Total network throughput is shown in the top figures, while inter-node throughput is depicted in the bottom ones..
These experimental results suggest that simply increasing intra-node network bandwidth does not guarantee improved performance. While multiple accelerators per end node are beneficial for minimizing latency during Tensor Parallelism (TP), they also increase the volume of inter-node traffic. If this traffic exceeds the capacity of the inter-node network or the NIC, performance degradation may occur even earlier when intra-node bandwidth is high.
4.4.3 Scale-Up conclusions
As the LLMs continue to grow, it becomes increasingly necessary to distribute model parameters across multiple accelerators, increasing the number of accelerators per end node. We analyzed the implications of scaling up the number of accelerators within an end node. We have observed that increasing the number of accelerators or the intra-node network speed can lead to bottlenecks at the destination NIC, because the intra-node network typically uses smaller payload sizes than the inter-node network, which results in a higher overhead due to conversion.
Therefore, accelerators’ scale-up within server nodes is effective where most communication occurs within the same node, such as in setups with intensive Tensor Parallelism (TP) within nodes and only Pipeline Parallelism (PP) or Data Parallelism (DP) across nodes. However, in scenarios where the memory available on a single node is insufficient to store a whole model, and TP must be applied across multiple nodes, inter-node communication becomes more frequent and critical. In such cases, system performance degrades significantly, and the benefits of scale-up diminish.
@@ 4.5 Scale-Out simulation experiments
In this section, we evaluate system performance as the number of end nodes increases, consequently increasing the total number of accelerators in the system. Table 3 summarizes the network configurations used in this scale-out evaluation. Each end node is equipped with $8$ accelerators. We evaluate two high-speed intra-node network configurations to assess their impact on overall performance. The rest of the network parameters are the same as in the previous experiments.
Table 3: Scale-Out network configurations.
@@
# Nodes Inter-node switches Accelerators Intra-node network speed
1 128 24 $1204$ 256 Gbps/accelerator
2 512 Gbps/accelerator
3 512 48 $4096$ 256 Gbps/accelerator
4 512 Gbps/accelerator
Figure 14 shows the network performance across the different configurations described in Table 3. When compared to the results in Section 4.4, we observe that the network behavior follows similar trends, with performance degradation occurring at comparable levels of traffic intensity.
(a) Configuration #1.
(b) Configuration #2.
(c) Configuration #3.
(d) Configuration #4.
Figure 14: Network performance as a function of traffic load (%) (network configurations from Table 2). Intra-node performance is shown in the upper section, while inter-node performance is depicted in the lower section.
For instance, in network configuration #2, which supports a theoretical peak throughput of $524288$ Gbps, the system achieves only: $\text{Throughput (Gbps)}=\frac{$16384$}{1.14\times 20}\times 128=$91980$~\text{Gbps}$ under the C1 traffic pattern. This corresponds to just 17.54% of the maximum network capacity. Similarly, in configuration #4, which has a theoretical peak throughput of $2097152$ Gbps, the observed throughput under C1 traffic is $367921$ Gbps—again, approximately 17.54% of peak capacity.
These results reinforce the observation that system scale-out can become critically limited by inter-node traffic patterns. When a significant portion of the traffic is directed between different end nodes, the link between the intra-node network and the NIC can become a severe bottleneck. Additionally, packetization overhead caused by differing payload sizes between intra- and inter-node communications can further degrade performance. As the number of end nodes increases, these limitations persist, making it essential to consider traffic patterns and hardware boundaries when scaling distributed systems for large-scale workloads such as LLM training and inference.
@@
[…]
Figure 15: Packet Latency divided into components versus traffic load for the traffic patterns C1-C4 and network configuration #4 from Table 2.
[…]
Figure 16: Network throughput vs traffic load (%) (network config. #4, Table 2).